### v2

# Thermodynamic bounds on energy use in Deep Neural Networks

Abstract
While Landauer’s principle sets a fundamental energy limit for irreversible digital computation, we show that Deep Neural Networks (DNNs) implemented on analog physical substrates can operate under markedly different thermodynamic constraints. We distinguish between two classes of analog systems: dynamic and quasi-static. In dynamic systems, energy dissipation arises from neuron resets, with a lower bound governed by Landauer’s principle. To analyse a quasi-static analog platform, we construct an explicit mapping of a generic feedforward DNN onto physical system described by a model Hamiltonian. In this framework, inference can proceed reversibly, with no minimum free energy cost imposed by thermodynamics. We further analyze the training process in quasi-static analog networks and derive a fundamental lower bound on its energy cost, rooted in the interplay between thermal and statistical noise. Our results suggest that while analog implementations can outperform digital ones during inference, the thermodynamic cost of training scales similarly in both paradigms.
The rapid progress in Artificial Intelligence (AI) has resulted in breakthrough applications across fields such as natural language processing [1, 2], computer vision [3, 4], and molecular biology [5]. As deep neural networks (DNNs) scale up in size and complexity [6, 2], the energy required for both training and inference is increasing rapidly [7], and it is projected to become a major contributor to overall energy consumption in the near future. In light of the need for energy-efficient DNNs, it is natural to explore the theoretical lower bounds on energy consumption for these systems.
In digital computing, Landauer’s principle [8, 9] provides a fundamental benchmark: erasing one bit of information costs at least $k_{B}T\ln 2$ in energy, reflecting the entropy reduction mandated by the Second Law of Thermodynamics. A naive application of Landauer’s limit to digital hardware suggests a minimal energy requirement of roughly $5\cdot 10^{-20}$ Joules per 16-bit floating point operation (FLOP). In practice, however, digital processors (e.g., the latest Nvidia GPU chips) operate at about $5\cdot 10^{-13}$ Joules per FLOP due to inefficiencies such as error correction, clocking, and other overheads. It is important to note that current digital implementations of neural network architectures are far from optimal in terms of energy use. In contrast, many analog platforms—including optical, electronic, quantum, and mechanical systems—potentially offer nearly reversible means for executing linear operations, which could dramatically reduce the energy cost associated with these computations [10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
Figure 1: A: Schematic representation of a DNN. B: A hypothetical analog implementation of DNN. A non-uniform linear medium may encode the linear transformations between the neuron layers without any energy loss, e.g., by elastic light scattering. The non-linear elements would only be used for implementing a neuron activation function, such as ReLU.
In a typical DNN, schematically shown in Figure 1a, computation proceeds in two steps: a linear transformation followed by a nonlinear activation [6, 20]:
$\mathbf{y}^{(n+1)}$ $=\widehat{\mathbf{W}}_{n}\mathbf{x}^{(n)}+\mathbf{b}^{(n)}$ (1)
$x_{i}$ $=f(y_{i})$ (2)
Here, $x_{i}$ denotes a real-valued variable assigned to neuron $i$, $\mathbf{x}^{(n)}$ is t he vector of neuron activities at layer $n$, and $f(y)$ is a nonlinear activation function. The weight matrix $\widehat{\mathbf{W}}_{n}$ and bias vector $\mathbf{b}^{(n)}$ specify the parameters of the layer.
Physical implementations of DNNs can be broadly categorized into three platforms: (i) digital computing, (ii) dynamic analog, and (iii) quasi-static analog systems. While the thermodynamic limits of digital computation are constrained by the Landauer principle, our discussion focuses on the latter two architectures.
In dynamic analog machines, physical signals propagate unidirectionally between consecutive layers. For concreteness, we consider a model system employing the ReLU (Rectified Linear Unit) activation function [20], which replaces Eq.(2) with
$x_{i}$ $=s_{i}y_{i}$ (3)
$s_{i}$ $=\Theta\left(y_{i}\right)$ (4)
where each neuron is assigned a state variable $s_{i}$, taking the values $s_{i}=1$ (active) or $s_{i}=0$ (inactive), and $\Theta(\cdot)$ denotes the Heaviside step function. Together, Eqs.(1) and (3) define a linear transformation that can be implemented via signal propagation in linear media—optical, electronic, mechanical, or others [10, 11, 12, 13, 14, 15, 18]. Importantly, there is no intrinsic thermodynamic bound on the energy cost of such linear transmission [9, 21]: while information capacity depends on the signal-to-noise ratio, the transmitted energy can, in principle, be fully recovered. Example include elastic light scattering in a non-uniform medium, illustrated in Figure 1b
By contrast, the neuron state variable $s_{i}$ is functionally equivalent to a physical bit. It can be switched from $s=0$ to $s=1$ by applying a signal above threshold ($y>0$), and its state can be read by probing its response to a small increment of the input signal $s_{i}=\partial x_{i}/\partial y_{i}$. At the end of each inference cycle, neurons are reset to the inactive state ($s_{i}=0$), a generic feature of many dynamic neuromorphic systems. According to Landauer’s principle—and more generally, the second law of thermodynamics—this reset operation incurs a minimal energy cost associated with entropy reduction:
$E_{\rm inf}>$ $-k_{B}T\nu\left(p\ln p+(1-p)\ln(1-p)\right)=$
$=k_{B}T\nu_{a}\left(1+\ln\frac{\nu}{\nu_{a}}\right)$ (5)
where $\nu_{a}$ is the average number of activated neurons per cycle, $\nu$ is the total neuron count, and $p=\nu_{a}/\nu\ll 1$ is the activation probability (assumed small). The standard Landauer result is recovered in the limit when each neuron is activated with probability 1/2 at each cycle, which gives $E_{\rm inf}>k_{B}T\nu\ln 2$.
The above result sheds new light on the long-standing challenge of quantifying the human brain’s computational power. Numerous studies, employing diverse methodologies, have produced estimates that vary by multiple orders of magnitude [22, 23, 24, 25]. The upper bound can be reliably set by Landauer’s limit: with the brain’s power consumption on the order of 10 W, this corresponds to an information processing rate of approximately $3\times 10^{21}$ bits per second. To set the lower bound, consider a hypothetical artificial neural network that matches the human brain’s neuron count — roughly $10^{11}$ neurons — with each neuron switching its activation state at a typical biological firing rate of about 10 Hz. As discussed above, much of this network’s functionality could, in principle, be implemented reversibly. Consequently, the minimum rate of irreversible operations, determined by the frequency of neuronal state transitions, may be as low as $10^{12}$ bit/s. Notably, this rate is within an order of magnitude of the information-theoretical bound calculated in Ref. [26]. While this reference DNN is unlikely to replicate the full performance of the actual brain, our estimate establishes a lower bound that complements the Landauer-based upper limit. Since both extremes are practically unattainable, it is reasonable to expect that the true computational power of the brain lies within a midrange window, approximately $10^{15}$ to $10^{19}$ bit/s. Indeed, multiple estimates tend to cluster within this range [25, 22, 23].
We now turn to quasi-static implementations of DNNs. Historically, much of neural network fied was shaped by analogies to statistical mechanics models such as the Sherrington–Kirkpatrick spin glass [27], which underpins Hopfield networks [6] and Boltzmann machines [28]. In these early models, inference corresponds to minimizing a Hamiltonian (or free energy at finite temperature), with neurons represented by binary spins coupled via symmetric interactions.
Subsequent developments introduced continuous variables $x_{i}$, nonlinear activation functions such as ReLU, and unidirectional couplings. This evolution enabled modern feedforward architectures and efficient training via backpropagation. However, unidirectional couplings are incompatible with the bidirectional interactions of the original Hamiltonians of the Boltzmann machine type, where couplings appear as terms like $J_{ij}x_{i}x_{j}$. Nevertheless, feedforward DNNs can still be exactly mapped onto a physical Hamiltonian of the following form:
$H=\sum_{i>i_{\rm in}}\frac{\kappa_{i}}{2}\left(x_{i}-f\left(\sum_{j}w_{ji}x_{j}+b_{i}\right)\right)^{2}$ (6)
Here, $w_{ji}=0$ for $j\geq i$, and $i_{\rm in}$ denotes the number of input neurons. For a given input ${\mathbf{x}}^{(0)}=\left(x_{1},\dots,x_{i_{\rm in}}\right)$, the Hamiltonian attains a trivial minimum of zero, corresponding to Eqs.(̃1)–(2) being satisfied.
This mapping allows one to recover many properties of classical Boltzmann machines. In particular, it enables the formulation of a finite-temperature version of the DNN, which introduces an additional Gaussian noise term into Eq.  (2): $x_{i}=f(y_{i})+\delta_{i}$, where $\langle\delta_{i}^{2}\rangle=k_{B}T/\kappa_{i}$. This noise propagates forward through the network according to Eqs.  (1)–(2). The corresponding free energy can be evaluated by Gaussian integration of the partition function near the minimum of the Hamiltonian, yielding $F=\frac{k_{B}T}{2}\sum_{i>i_{\rm in}}\ln\left(\frac{\kappa_{i}}{k_{B}T}\right)$. Importantly, this free energy is independent of the model parameters $w_{ij}$, $b_{i}$, and the input values. It can therefore be regarded as a constant, much like the kinetic energy contribution that has been omitted from the Hamiltonian in Eq.(6). A key implication of this construction is that inference in a quasi-static network can, in principle, be performed in a thermodynamically reversible manner, without any global entropy production, $\Delta S=0$. In other words: Thermodynamics imposes no lower bound on the energy cost of quasi-static inference:
$E^{\min}_{\rm inf}=0$ (7)
To operate in this reversible regime, the system must remain at constant temperature, with input changes occurring slowly relative to the relaxation times of all internal variables. An important distinction from digital computing arises here: the quasi-static system described above has a single free energy minimum and does not experience ergodicity breaking. In contrast, each physical bit in a digital computer is implemented as a pair of (meta)stable states with lifetimes exceeding a single computational cycle. This multiplicity of stable states ultimately gives rise to the Landauer bound on minimal energy dissipation, even in the quasi-static limit.
We now proceed to discuss the thermodynamic bounds on energy use during the training of DNNs. This problem has been addressed in the past, primarily from information-theoretical point of view. In particular, in Refs. [29, 30] the mutual information between true and inferred values was shown to set the lower bound on the free energy cost of training. This bound however does not take into account the actual complexity of the underlying network, and is likely to be overoptimistic. For dynamic (neuromorphic) machines, a generic analysis is complicated due to the wide range of plausible physical implementations and training rules. Thus, this work limits its scope to discussing the training of quasi-static DNNs. Fortunately, the same physical model as above, Eq. (6), naturally describes the learning process. In a standard setup, training aims to minimize certain loss functions, such as the mean square error (MSE), employing stochastic gradient descent through error backpropagation. Since physical relaxation processes inspired these techniques, it should not be surprising that the learning procedure can be realized as an actual physical process. Indeed, this has been demonstrated in various model physical systems.
To train the physical DNN described by Hamiltonian Eq. (6), we do not introduce any additional loss function. Instead, each entry in the training dataset constrains both input and output variables to their respective values: $({\mathbf{x}}^{(0)},{\mathbf{x}}^{(h)})_{\alpha}$. Here index $\alpha$ enumerates training data entries, and $h$ denotes the DNN depth, i.e., the index of its output layer. We assume a significant separation between two relaxation time constants: (i) the minimum inference time $\tau_{inf}$, set by relaxation of the neuron variables $x_{i}$, and (ii) the training time $\tau_{train}$, associated with the slow annealing of the model parameters - weights $w_{ij}$ and biases $b_{i}$. Since both inputs and outputs are constrained during training, the Hamiltonian generally cannot achieve its trivial minimum $H=0$ for fixed model parameters. Physically, it results in the network being strained and non-zero gradients in the parameter space emerging:
$\frac{\partial H}{\partial b_{i}}$ $\equiv\sigma_{i}=\kappa_{i}\left(f(y_{i})-x_{i}\right)f^{\prime}(y_{i})$ (8)
$\frac{\partial H}{\partial w_{ji}}$ $=\sigma_{i}x_{j}$ (9)
Here $y_{i}=\sum_{j}w_{ji}x_{j}+b_{i}$. Values of $\sigma_{i}$ that can be interpreted as local stress in the network, are obtained by minimizing $H$ with respect to $x_{i}$:
$\sigma_{i}=f^{\prime}(y_{i})\sum_{j}w_{ij}\sigma_{j}$ (10)
Backpropagation naturally emerges in the system: the stress value in the output layer is proportional to the error, $\sigma_{i}=\kappa_{i}(x_{i}-x_{i,\alpha})f^{\prime}(y_{i})$, and can be computed recursively across the network by moving backward, layer-by-layer. The calculated derivatives of the physical Hamiltonian are proportional to those of a conventional MSE loss function employed in the standard DNN training procedure. Thus, the stochastic gradient descent can be directly implemented through physical annealing of the parameters. This is quite natural, as our discussion here essentially parallels the classical approach to learning in a Bolzmann Machine [28]. It also echoes many approaches to in-situ physical learning proposed and implemented in recent year [12, 13, 14, 15, 16, 17, 18, 19]. It should be emphasized that there are multiple ways of mapping DNN onto a physical Hamiltonian. However, in order for the correct learning rules to emerge from physical dynamics (i) there must be clear separation of the two relaxation times, $\tau{\rm inf}$ and $\tau_{\rm train}$, and (ii) the minimum free energy at the inference phase should be independent of model parameters. This ensures that the annealing of the physical system in the parameter space will indeed lead to relaxation of stresses, and minimization of error.
We consider the DNN being sequentially exposed to the training dataset, i.e. input/output pairs $({\mathbf{x}}^{(0)},{\mathbf{x}}^{(h)})_{\alpha}$, $\alpha=1,...,D$. In the standard in-silico training, data entries are typically grouped into batches, and the average gradient over the batch is calculated. The finite size of a batch results in a variation of the gradient from its average value across the entire dataset, giving rise to an effective stochastic noise. In physical training, the batching appears naturally, due to the separation of the time scales: inference time $\tau_{\rm inf}$ and training time $\tau_{\rm train}$. The data entries used within a single inference time window (i.e. within the relaxation time of neuron variables $x_{i}$) constitute a single data batch, as the gradient $\nabla_{\theta}H_{\alpha}$ in the parameter space is effectively averaged over them. In other words, the ratio of time constants $\tau_{\rm train}/\tau_{\rm inf}$ corresponds to the effective number of training epochs.
One can now describe the annealing of parameters by a set of standard Langevin equations:
$\dot{\theta}_{k}=-\mu_{k}\partial_{k}F(t)+\eta_{k}(t)$ (11)
$\langle\eta_{k}(t)\eta_{k}(t^{\prime})\rangle=2k_{B}T\mu_{k}\delta(t-t^{\prime})$ (12)
Here $\theta_{k}$ is a model ($w_{ij}$, or $b_{i}$), and $\partial_{k}F_{\alpha}=\langle H_{\alpha}\rangle/\partial\theta_{k}$ is the corresponding derivative of the free energy $F_{\alpha}$, for a training data entry $({\mathbf{x}}^{(0)},{\mathbf{x}}^{(h)})_{\alpha}$. $\eta_{k}(t)$ is the Gaussian thermal noise that satisfies the Fluctuation Dissipative Theorem Let $\overline{\partial_{k}F}$ be a value of the derivative time-averaged over timescale between $\tau_{\rm inf}$ and $\tau_{\rm train}$, i.e., over multiple entries. By separating this slow part of $\partial_{k}F(t)$ from its entry-to-entry variation, one can rewrite the above Langevin equation as:
$\dot{\theta}_{k}=-\mu_{k}\overline{\partial_{k}F}+\xi_{k}(t)+\eta_{k}(t)$ (13)
$\langle\xi_{k}(t)\xi_{k}(t^{\prime})\rangle=\frac{\mu_{k}^{2}\tau_{\rm train}}{D}{\rm var}(\partial_{k}F)\delta(t-t^{\prime})$ (14)
Here, the magnitude of the statistical noise $\xi(t)$ was obtained by assuming the consecutive data entries to be mutually independent. ${\rm var}(\partial_{k}F)$ refers to the variance of the corresponding derivative in the dataset, and $\tau_{\rm train}/D$ is the exposure time per entry. Remarkably, this noise level is closely related to the total unrecoverable work done due to dissipation during the training, which follows from Eq. (13):
$W=\int_{0}^{\tau_{\rm train}}\sum_{k}\dot{\theta}_{k}\partial_{k}F(t)dt\approx\tau_{\rm train}\mu\langle\left|\nabla_{\theta}F\right|^{2}\rangle$ (15)
We are now in a position to estimate the thermodynamic lower bound on the energy required to train a quasi-static analog network. Let us assume that we have already identified the optimal regime for *in silico* DNN training. As discussed above, this regime can be directly implemented through annealing of the physical system described by the Hamiltonian in Eq. (6), using an appropriately chosen ratio of time constants, $\tau_{\rm train}/\tau_{\rm inf}$. In this setting, the system follows the stochastic gradient descent dynamics given by Eq. (13), but without thermal noise $\eta(t)$.
When thermal effects are introduced, one expects training performance to eventually degrade; however, this degradation only becomes significant once the magnitude of the thermal noise $\eta(t)$ becomes comparable to that of the statistical noise $\xi$. By comparing Eqs. (12) and (14), and using the expression for work done, Eq. (15), we arrive at the following lower bound for the total energy required for training, in terms of dataset size $D$ and number of model parameters $N$:
$E_{\rm train}\gtrsim 2NDk_{B}T\frac{\langle\left|\nabla_{\theta}F\right|^{2}\rangle}{{\rm var}\left(\nabla_{\theta}F\right)}\approx 2NDk_{B}T.$ (16)
This result is both simple and remarkable. It bares a strong similarity to the Thermodynamic Uncertainty Relationship (TUR), that has recently gained prominence in the context of non-equilibrium statistical mechanics [31]. However, in TUR, the expected values and variances refer to fluctuations of a non-equilibrium system, while in the current context, they emerge from the statistics of the training dataset.
Perhaps unexpectedly, the Eq. (16) is not too different from the well-known estimates for the computational cost of in silico DNN training. Particularly large language models (LLMs) and transformer architectures require approximately $6ND$ floating point operations (FLOPs) [2]. If we assume 16-bit precision per FLOP and apply Landauer’s principle, that estimate translates to a minimal digital training energy of:
$E_{\rm train}^{(\rm dig)}\gtrsim 10^{2}N_{a}Dk_{B}T$ (17)
Note that in modern Mixture of Experts (MoE) architectures, only a small subset of network parameters is activated for each training sample. Thus, the effective number of active parameters $N_{a}$ may be much smaller than the total parameter count $N$ appearing in Eq. (16) for analog training. This suggests that, at least in principle, in silico training may outperform analog training, in sharp contrast to inference, where dynamic and quasi-static analog systems can be vastly more energy efficient than digital ones, as established by Eqs.(5) and (7).
Furthermore, digital systems offer an additional advantage: once trained, neural network models can be copied and deployed essentially for free—an operation that is highly nontrivial for physical systems trained via slow annealing. That being said, present-day digital computers still operate at least 7 orders of magnitude above Landauer’s limit, with no clear pathway for dramatically closing this gap. In contrast, the physical realization discussed in this work offers a plausible route to building systems capable of operating near the thermodynamic bounds on energy efficiency. In Table 1 we present a comparison between our results, Eqs. (5),(7) (16), and the actual energy use of the modern LLMs. We also include the respective estimates of minimal energy use by a digital computer, set by the Landauer limit.
Function Energy Use ($J/token$)
(model) Current     Landauer limit Analog bounds
Inference
(Llama 65B)
$4$ $5\cdot 10^{-8}$
$10^{-14}$ (dynamic)
00 (quasi-static)
Training
(DeepSeek V3 )
.2.2.2.2 $5\cdot 10^{-8}$ $5\cdot 10^{-9}$
Table 1: Comparison of our results with actual energy use by modern LLMs: LLama 65B (inference), and DeepSick V3 (training), as well as with the respective Landauer limits. The estimates are based on data from Refs. [32, 33], as well as official specifications of Nvidia GPUs A100 and H800. Analog bounds are given by Eqs. (5),(7) and (16).
Since the thermodynamic bounds determined in this work, Eq. (5) and (16) scale with $T$, one might anticipate that lowering the operating temperature would reduce the energy cost. While operating at a reduced temperature $T^{*}<T$ may offer practical benefits, it does not circumvent the fundamental limitations imposed by the Second Law of Thermodynamics. To demonstrate this, consider a physical DNN maintained at a temperature $T^{*}$ below the ambient temperature $T$. The energy $E$ used in operation must eventually be removed as heat. According to the Second Law, this removal requires performing work $W$ such that the overall entropy does not decrease:
$\frac{W+E}{T}-\frac{E}{T^{*}}\geq 0$ (18)
From this inequality, one obtains
$E_{\min}(T)=W_{\min}+E_{\min}(T^{*})=\frac{T}{T^{*}}E_{\min}(T^{*}),$ (19)
demonstrating that cooling cannot beat the thermodynamic energy bounds. Furthermore, the physical training Eq. (11) should typically be performed at a higher temperature than the inference to allow the parameter annealing over the training time, and preventing it during the normal operations.
Finally, it is important to recognize that the conventional use of energy or work as thermodynamic ”currencies” is largely a legacy of the Industrial Revolution. Throughout this paper, where we referred to ”energy cost,” a more precise formulation would involve the appropriate version of free energy. In the context of the information age, however, a more natural and general measure of thermodynamic cost is negentropy—a concept promoted by Schrödinger in his book What is Life? [34] and later formalized by L. Brillouin [35, 36]. Negentropy can be defined as:
$J\equiv-\frac{\Delta S}{k_{B}\ln 2}=\frac{F}{k_{B}T\ln 2},$ (20)
where $F$ is the Helmholtz free energy of the system of interest, and $S$ is the total entropy, including the thermal bath. We have rescaled negentropy to ensure that it is measured in bits, same units Shannon’s information entropy $H$. Brillouin reinterpreted the Second Law of Thermodynamics as Information Principle, nearly a decade before Landauer’s work:
$\Delta(J+H)\leq 0,$ (21)
In the context of the present study, our key results, Eqs. (5), (7), and (16), should be reformulated as lower bounds on the negentropy required for inference and training in DNNs.
In conclusion, we have investigated fundamental thermodynamic bounds governing energy consumption in DNNs, during both inference and training phases. For dynamic analog systems, which can be broadly classified as neuromorphic, we concluded that the thermopdynamic bound on inference energy is primarily determined by the need to reset the neuron state after each cycle, leading to Eq. (5). We also demonstrated that quasi-static analog DNNs, described through an explicit Hamiltonian, can theoretically achieve reversible inference operations with no theoretical minimum of energy use imposed by thermodynamics. Training in quasi-static analog systems, however, is fundamentally constrained by a different thermodynamic bound, Eq. (16). This bound has the same scaling relationship with training dataset size $D$ and parameter number $N$ as the computing power needed for in silico training. The digital platforms retain certain pragmatic advantages, including the ability to optimize the training protocol by using MoE architecture, and the ease and negligible energy cost of duplicating trained models. Nevertheless, our analysis demonstrates that analog implementations hold significant promise for surpassing current digital systems in terms of practical energy efficiency.
Acknowledgements.
Acknowledgments. This research was done at and used resources of the Center for Functional Nanomaterials, which is a U.S. DOE Office of Science User Facility, at Brookhaven National Laboratory under Contract No. DE-SC0012704.