### v3

# Thermodynamic bounds on energy use in quasi-static Deep Neural Networks

Abstract
The rapid growth of deep neural networks (DNNs) has brought increasing attention to their energy use during training and inference. Here we establish the thermodynamic bounds on energy consumption in quasi-static analog DNNs by mapping modern feedforward architectures onto a physical free-energy functional. This framework provides a direct statistical-mechanical interpretation of quasi-static DNNs. As a result, inference can proceed in a thermodynamically reversible manner, with vanishing minimal energy cost, in contrast to the Landauer limit that constrains digital hardware. Importantly, inference corresponds to relaxation to a unique free-energy minimum with $F_{\min}=0$, allowing all constraints to be satisfied without residual stress. By comparison, training overconstrains the system: simultaneous clamping of inputs and outputs generates stresses that propagate backward through the architecture, reproducing the rules of backpropagation. Parameter annealing then relaxes these stresses, providing a purely physical route to learning without an explicit loss function. We further derive a universal lower bound on training energy, $E_{\text{train}}\gtrsim 2NDk_{B}T$, which scales with both the number of trainable parameters and the dataset size.
I Introduction
The rapid progress in Artificial Intelligence (AI) has resulted in breakthrough applications across fields such as natural language processing [1, 2], computer vision [3, 4], and molecular biology [5]. As deep neural networks (DNNs) scale up in size and complexity [6, 2], the energy required for both training and inference is increasing rapidly [7], and it is projected to become a major contributor to overall energy consumption in the near future. In light of the need for energy-efficient DNNs, it is natural to explore the theoretical lower bounds on energy consumption for these systems.
In digital computing, Landauer’s principle [8, 9] provides a fundamental benchmark: erasing one bit of information costs at least $k_{B}T\ln 2$ in energy, reflecting the entropy reduction mandated by the Second Law of Thermodynamics. A naive application of Landauer’s limit to digital hardware suggests a minimal energy requirement of roughly $5\cdot 10^{-20}$ Joules per 16-bit floating point operation (FLOP). In practice, however, digital processors (e.g., the latest Nvidia GPU chips) operate at about $5\cdot 10^{-13}$ Joules per FLOP due to inefficiencies such as error correction, clocking, and other overheads.
It is important to note that current digital implementations of neural network architectures are far from optimal in terms of energy use. Alternative analog platforms, including optical, electronic, quantum, and mechanical systems, could dramatically reduce the energy cost associated with these computations [10, 11, 12, 13, 14, 15, 16, 17, 18, 19].The physical implementations of DNNs can be broadly categorized into three platforms: (i) digital computing, (ii) dynamic analog (neuromorphic), and (iii) quasi-static analog systems. In this paper, we address the problem of the thermodynamic bound on energy required for inference and training of DNNs, focusing on the quasi-static case.
Figure 1: Schematic representation of a Deep Neural Network.
II Inference energy in quasi-static DNNs
In a typical DNN, schematically shown in Figure 1, computation proceeds in two steps: a linear transformation followed by a nonlinear activation [6, 20]:
$\mathbf{y}^{(n+1)}$ $=\widehat{\mathbf{W}}_{n}\mathbf{x}^{(n)}+\mathbf{b}^{(n)}$ (1)
$x_{i}$ $=f(y_{i})$ (2)
Here, $x_{i}$ denotes a real-valued variable assigned to neuron $i$, $\mathbf{x}^{(n)}$ is t he vector of neuron activities at layer $n$, and $f(y)$ is a nonlinear activation function. The weight matrix $\widehat{\mathbf{W}}_{n}$ and bias vector $\mathbf{b}^{(n)}$ specify the parameters of the layer.
Historically, much of the neural network field was shaped by analogies to statistical mechanics models such as the Sherrington–Kirkpatrick spin glass [21], which underpins Hopfield networks [6] and Boltzmann machines [22]. In these early models, inference corresponds to minimizing a Hamiltonian (or free energy at finite temperature), with neurons represented by binary spins coupled via symmetric interactions.
Subsequent developments introduced continuous variables $x_{i}$, nonlinear activation functions such as ReLU, and unidirectional couplings. This evolution enabled modern feedforward architectures and efficient training via backpropagation. However, unidirectional couplings are incompatible with the bidirectional interactions of the original Hamiltonians of the Boltzmann machine type, where couplings appear as terms like $J_{ij}x_{i}x_{j}$. Nevertheless, feedforward DNNs can still be exactly mapped onto a physical free energy of the following form:
$F=\sum_{i>i_{\rm in}}\frac{\kappa_{i}}{2}\left(x_{i}-f\left(\sum_{j}w_{ji}x_{j}+b_{i}\right)\right)^{2}$ (3)
Here, $w_{ji}=0$ for $j\geq i$, and $i_{\rm in}$ denotes the number of input neurons. For a given input ${\mathbf{x}}^{(0)}=\left(x_{1},\dots,x_{i_{\rm in}}\right)$, the free energy attains a trivial minimum of zero, corresponding to Eqs.(̃1)–(2) being satisfied.
This mapping allows one to recover many properties of classical Boltzmann machines. In particular, it enables the formulation of a finite-temperature version of the DNN, which introduces an additional Gaussian noise term into Eq.  (2): $x_{i}=f(y_{i})+\delta_{i}$, where $\langle\delta_{i}^{2}\rangle=k_{B}T/\kappa_{i}$. This noise propagates forward through the network according to Eqs.  (1)–(2). The fluctuational corrections to the free energy can be evaluated by Gaussian integration of the partition function near the minimum of Eq.(3), $F=0$, yielding
$F=\frac{k_{B}T}{2}\sum_{i>i_{\rm in}}\ln\left(\frac{\kappa_{i}}{k_{B}T}\right)$ (4)
Importantly, this free energy is independent of the model parameters $w_{ij}$, $b_{i}$, and the input values. It can therefore be regarded as a constant, much like the kinetic energy contribution that has been omitted from Eq.(3). A key implication of this construction is that inference in a quasi-static network can, in principle, be performed in a thermodynamically reversible manner, without any global entropy production, $\Delta S=0$. In other words: Thermodynamics imposes no lower bound on the energy cost of quasi-static inference:
$E^{\min}_{\rm inf}=0$ (5)
To operate in this reversible regime, the system must remain at constant temperature, with input changes occurring slowly relative to the relaxation times of all internal variables. An important distinction from digital computing arises here: the quasi-static system described above has a single free energy minimum and does not experience ergodicity breaking. In contrast, each physical bit in a digital computer is implemented as a pair of (meta)stable states with lifetimes exceeding a single computational cycle. This multiplicity of stable states ultimately gives rise to the Landauer bound on minimal energy dissipation, even in the quasi-static limit.
III Training through parameter annealing
We now proceed to discuss the thermodynamic bounds on energy use during the training of DNNs. This problem has been addressed in the past, primarily from an information-theoretical point of view. In particular, in Refs. [23, 24] the mutual information between true and inferred values was shown to set the lower bound on the free energy cost of training. This bound however does not take into account the actual complexity of the underlying network, and is likely to be overoptimistic.
Remarkably, the same physical model as above, Eq. (3), naturally describes the learning process. In a standard setup, training aims to minimize certain loss functions, such as the mean square error (MSE), employing stochastic gradient descent through error backpropagation. Since physical relaxation processes inspired these techniques, it should not be surprising that the learning procedure can be realized as an actual physical process. Indeed, this has been demonstrated in various model physical systems. To train the physical DNN described by free energy, Eq. (3), we do not introduce any additional loss function. Instead, each entry in the training dataset constrains both input and output variables to their respective values: $({\mathbf{x}}^{(0)},{\mathbf{x}}^{(h)})_{\alpha}$. Here index $\alpha$ enumerates training data entries, and $h$ denotes the DNN depth, i.e., the index of its output layer. While this approach is analogous to the use of clumped and free states in the context physical learning, important distinction is that our learning rule is given by regular relaxation of parameters towards the Free energy minimum. This is possible because the inference corresponds to a minimum at $F=0$, and thus free energy itself acts as a loss function for overconsiderate (clumped) state. We assume a significant separation between two relaxation time constants: (i) the minimum inference time $\tau_{inf}$, set by relaxation of the neuron variables $x_{i}$, and (ii) the training time $\tau_{train}$, associated with the slow annealing of the model parameters - weights $w_{ij}$ and biases $b_{i}$. Since both inputs and outputs are constrained during training, the free energy generally cannot achieve its trivial minimum $H=0$ for fixed model parameters. Physically, it results in the network being strained and non-zero gradients in the parameter space emerging:
$\frac{\partial F}{\partial b_{i}}$ $\equiv\sigma_{i}=\kappa_{i}\left(f(y_{i})-x_{i}\right)f^{\prime}(y_{i})$ (6)
$\frac{\partial F}{\partial w_{ji}}$ $=\sigma_{i}x_{j}$ (7)
Here $y_{i}=\sum_{j}w_{ji}x_{j}+b_{i}$. Values of $\sigma_{i}$ that can be interpreted as local stress in the network, are obtained by minimizing $F$ with respect to $x_{i}$:
$\sigma_{i}=f^{\prime}(y_{i})\sum_{j}w_{ij}\sigma_{j}$ (8)
Backpropagation naturally emerges in the system: the stress value in the output layer is proportional to the error, $\sigma_{i}=\kappa_{i}(x_{i}-x_{i,\alpha})f^{\prime}(y_{i})$, and can be computed recursively across the network by moving backward, layer-by-layer. The calculated derivatives of the physical free energy are proportional to those of a conventional MSE loss function employed in the standard DNN training procedure. Thus, the stochastic gradient descent can be directly implemented through physical annealing of the parameters. This is quite natural, as our discussion here essentially parallels the classical approach to learning in a Boltzmann Machine [22]. It also echoes many approaches to in-situ physical learning proposed and implemented in recent years [12, 13, 14, 15, 16, 17, 18, 19]. It should be emphasized that there are multiple ways of mapping DNN onto a physical free energy. However, in order for the correct learning rules to emerge from physical dynamics (i) there must be a clear separation of the two relaxation times, $\tau_{\rm inf}$ and $\tau_{\rm train}$, and (ii) the minimum free energy at the inference phase should be independent of model parameters. This ensures that the annealing of the physical system in the parameter space will indeed lead to relaxation of stresses, and minimization of error.
IV Thermodynamic bound on training energy
We consider sequential exposure to a dataset of $D$ input/output pairs $({\mathbf{x}}^{(0)},{\mathbf{x}}^{(h)})_{\alpha}$, $\alpha=1,\dots,D$. In in-silico training, one averages gradients over mini-batches; in the our physical setting, an analogous averaging arises naturally from the separation of timescales: neuron variables relax on a fast timescale $\tau_{\rm inf}$, while parameters evolve on a slow timescale $\tau_{\rm train}$. Over an interval of duration $\tau_{\rm inf}$, the effective driving gradient is thus a finite-sample average, which fluctuates around the dataset mean. Note that in a general case, exposing the system to multiple data entries within $\tau_{\rm inf}$ would lead to a violation of the quasi-staticity. For this reason, below we focus on the case when the switch between consecutive entries is slow enough, i.e., occurs on the timescale of at least $\tau_{\rm inf}$.
Let $\theta_{k}\in\{w_{ij},b_{i}\}$ denote a parameter with mobility $\mu_{k}$. Its dynamics is modeled as
$\dot{\theta}_{k}$ $=-\mu_{k}\,\partial_{k}F(t)+\eta_{k}(t),$ (9)
$\langle\eta_{k}(t)\eta_{k}(t^{\prime})\rangle$ $=2k_{B}T\,\mu_{k}\,\delta(t-t^{\prime}),$ (10)
where $\eta_{k}$ is thermal noise and $\partial_{k}F(t)$ is the free-energy gradient. We decompose this gradient into a slow mean and a fast fluctuation,
$\color[rgb]{0,0,0}\partial_{k}F(t)=\overline{\partial_{k}F}+\delta(\partial_{k}F)$}. (11)
The fast part originates from finite-sample averaging over intervals $\sim\tau_{\rm inf}$ and acts as an additional effective noise, which we denote $\xi_{k}(t)$. Under the assumptions of (i) independence between successive samples and (ii) a clear timescale separation $\tau_{\rm inf}\ll\tau_{\rm train}$ so that central-limit arguments apply, one obtains
$\dot{\theta}_{k}$ $=-\mu_{k}\,\overline{\partial_{k}F}+\xi_{k}(t)+\eta_{k}(t)$ (12)
$\langle\xi_{k}(t)\xi_{k}(t^{\prime})\rangle$ $=\frac{\mu_{k}^{2}\,\tau_{\rm train}}{D}\;{\rm var}(\partial_{k}F)\,\delta(t-t^{\prime})$ (13)
The factor $\tau_{\rm train}/D$ is the average exposure time per data entry. Note that $\langle.\rangle$ represents a combination of two types of averaging. One of them is the conventional thermal average, the other is the averaging over the stochastic ensemble of possible training sets.
We next relate dissipation to parameter drift. The power injected into the system by changing $\theta$ is given by standard thermodynamic expression
$\color[rgb]{0,0,0}\dot{W}(t)=\sum_{k}\dot{\theta}_{k}\,\partial_{k}F(t).$} (14)
Integrating over the training interval gives the total work,
$\color[rgb]{0,0,0}\langle W\rangle=\int_{0}^{\tau_{\rm train}}\!\!\sum_{k}\mu_{k}\langle(\partial_{k}F(t))^{2}\rangle dt\geq\tau_{\rm train}\sum_{k}\mu_{k}{\rm var}(\partial_{k}F)$} (15)
Here, we used the fact that $\langle(\partial_{k}F(t))^{2}\rangle\geq{\rm var}(\partial_{k}F)$. In other words, we dropped the part of the work associated with relaxation along the slow component of the force. The time-averaged value of ${\rm var}(\partial_{k}F)$ during the training process has been replaced with its ensemble average, due to the independence of successive data entries.
In order to finalize our derivation of the training energy bound, we note that for $\langle\xi^{2}\rangle\gg\langle\eta^{2}\rangle$ the gradient signal extracted from data dominates the overall dynamics and learning proceeds reliably. Conversely, thermal noise start influencing the dynamics once $\langle\eta^{2}\rangle\simeq\langle\xi^{2}\rangle$. If we assume the combination of $\tau_{inf}$ and $\tau_{train}$ to be chosen to optimize the performance of the model, a noticeable increase in stochasticity will result in its deterioration. This noise balance criterion sets the maximum strength of thermal noise that the models would tolerate. Based on Eqs. (10) and (13), it gives the following bound for the optimal operating temperature:
$\color[rgb]{0,0,0}\frac{\mu_{k}^{2}\,\tau_{\rm train}}{D}\;{\rm var}(\partial_{k}F)\gtrsim 2k_{B}T\,\mu_{k}$} (16)
After using Eq. (15), and summing over $N$ parameters, this yields a lower bound for the total training energy cost,
$\color[rgb]{0,0,0}E_{\rm train}=\langle W\rangle\;\gtrsim\;2\,N\,D\,k_{B}T.$} (17)
Equation (17) shows that even in an ideal quasi-static analog implementation, reliable training requires finite dissipation scaling with the number of parameters $N$ and dataset size $D$. This result is both simple and remarkable. It bears a strong similarity to the Thermodynamic Uncertainty Relationship (TUR), which has recently gained prominence in the context of non-equilibrium statistical mechanics [25]. However, in TUR, the expected values and variances refer to fluctuations of a non-equilibrium system, while in the current context, they emerge from the statistics of the training dataset.
V Discussion
Perhaps unexpectedly, the Eq. (17) is not too different from the well-known estimates for the computational cost of in silico DNN training. Particularly large language models (LLMs) and transformer architectures require approximately $6ND$ floating point operations (FLOPs) [2]. If we assume 16-bit precision per FLOP and apply Landauer’s principle, that estimate translates to a minimal digital training energy of:
$E_{\rm train}^{(\rm dig)}\gtrsim 10^{2}N_{a}Dk_{B}T$ (18)
Note that in modern Mixture of Experts (MoE) architectures, only a small subset of network parameters is activated for each training sample. Thus, the effective number of active parameters $N_{a}$ may be much smaller than the total parameter count $N$ appearing in Eq. (17) for analog training. This suggests that, at least in principle, in silico training may outperform analog training, in sharp contrast to inference, where dynamic and quasi-static analog systems can be vastly more energy efficient than digital ones, as established by Eq. (5).
Furthermore, digital systems offer an additional advantage: once trained, neural network models can be copied and deployed essentially for free—an operation that is highly nontrivial for physical systems trained via slow annealing. That being said, present-day digital computers still operate at least 7 orders of magnitude above Landauer’s limit, with no clear pathway for dramatically closing this gap. In contrast, the physical realization discussed in this work may provide a plausible route to building systems capable of operating near the thermodynamic bounds. In Table 1 we present a comparison between our results, Eqs. (5),(17), and the actual energy use of the modern LLMs. We also include the respective estimates of minimal energy use by a digital computer, set by the Landauer limit.
Function Energy Use ($J/token$)
(model) Current     Landauer limit Analog bound
Inference
(Llama 65B)
44 $5\cdot 10^{-8}$ 0
Training
(DeepSeek V3 )
.2.2 $5\cdot 10^{-8}$ $5\cdot 10^{-9}$
Table 1: Comparison of our results with actual energy use by modern LLMs: LLama 65B (inference), and DeepSick V3 (training), as well as with the respective Landauer limits. The estimates are based on data from Refs. [26, 27], as well as official specifications of Nvidia GPUs A100 and H800. Analog bounds are given by Eqs. (5) and (17).
Figure 2: Illustration of the impossibility of beating the thermodynamic bound by cooling the physical DNN.
Since the thermodynamic bound on training energy (17) scales with $T$, one might anticipate that lowering the operating temperature would reduce the energy cost. While operating at a reduced temperature $T^{*}<T$ may offer practical benefits, it does not circumvent the fundamental limitations imposed by the Second Law of Thermodynamics. To demonstrate this, consider a physical DNN maintained at a temperature $T^{*}$ below the ambient temperature $T$. The energy $E$ used in operation must eventually be removed as heat. According to the Second Law, this removal requires performing work $W$ such that the overall entropy does not decrease:
$\frac{W+E}{T}-\frac{E}{T^{*}}\geq 0$ (19)
From this inequality, one obtains
$E_{\min}(T)=W_{\min}+E_{\min}(T^{*})=\frac{T}{T^{*}}E_{\min}(T^{*}),$ (20)
demonstrating that cooling cannot beat the thermodynamic energy bounds. Furthermore, the physical training Eq. (9) should typically be performed at a higher temperature than the inference to allow the parameter annealing over the training time, and preventing it during the normal operations.
VI Conclusions
We have developed a thermodynamic framework for quasi-static analog DNNs by mapping them onto a physical system described by free energy $F$, Eq. (3). This approach provides a direct statistical mechanical interpretation of neural computation. Inference in such systems corresponds to quasi-static relaxation of neuron variables to a unique global minimum of the free energy. As a result, inference can proceed in a thermodynamically reversible manner, with no entropy production and zero minimal energy cost, Eq. 5). This radically distinguishes the quasi-static analog platform from digital hardware, where the nessesary multistability of bits leads to Landauer’s limit.
Importantly, for any fixed input the minimum satisfies $F_{\min}=0$ (up to a constant fluctuational correction, Eq.(4)), which means all constraints can be satisfied simultaneously without residual stress. This observation places quasi-static inference in close analogy with isostatic elastic networks, where the number of constraints matches the available degrees of freedom, allowing the system to accommodate them without storing elastic energy. By contrast, the training overconstrains the system, as both inputs and target outputs are clamped simultaneously. In this case, the network cannot reach $F=0$; instead, finite stresses $\sigma_{i}$ appear in the free-energy landscape [Eqs. (5)–(7)]. These stresses propagate backward through the architecture and serve as physical analogs of the error signals in backpropagation. When parameters such as weights and biases evolve slowly compared to neuron relaxation, their annealing naturally relaxes stresses, thereby reducing error. In this way, training arises intrinsically from the physics of the system, without the need for an explicit externally defined loss function. This feature has profound implications for the physical learning community: it demonstrates that learning in analog substrates can be realized through pure parameter annealing, guided only by thermodynamic relaxation.
At the same time, our analysis shows that such training is irreducibly dissipative. The lower bound on training energy, Eq. (17), scales extensively with both the number of parameters $N$ and the dataset size $D$. This identifies a fundamental asymmetry: inference can, in principle, be reversible, but adaptation through learning necessarily incurs finite thermodynamic cost. The structure of Eq. (17) also draws a striking parallel with thermodynamic uncertainty relations, extending their scope to learning in artificial neural systems.
In summary, the free-energy model provides a unifying statistical-mechanical foundation for DNNs, clarifying when energy-efficient analog computation is possible and when thermodynamics imposes unavoidable limits. It highlights quasi-static analog inference as a plausible route toward reversible computing, while also establishing that training, although elegantly realized through intrinsic parameter annealing, carries a fundamental energetic price. For the design of future physical learning machines, these insights suggest a path toward architectures where computation and adaptation emerge not from imposed algorithms, but from the thermodynamics of relaxation itself.
Acknowledgements.
Acknowledgments. This research was done at and used resources of the Center for Functional Nanomaterials, which is a U.S. DOE Office of Science User Facility, at Brookhaven National Laboratory under Contract No. DE-SC0012704.